# Rotational Taylor dispersion in linear flows

### v1

@@ 3.2 Extensional flow
As another case study, we now consider an extensional flow where the angular velocity is $\bm{\Omega}=Pe\,B\cos(2\phi)\bm{e}_{z}$. The extensional flow tends to align the particle orientation with the extensional axis (see figure 5). The particle has a nonzero angular velocity when the orientation deviates from the extensional axis. As shown in figure 5(b), the direction of rotation depends on the particle orientation relative to the extensional axis.
@@
Figure 6: Plots of the effective dispersion coefficient as a function of $Pe\,B$ in the extensional flow. Circles are results obtained from the numerical solution of the Taylor dispersion theory, squares are results calculated from BD simulations, and the dashed line denotes the leading-order asymptotic solution for small $Pe\,B$ given in (40).
In figure 6, we plot the effective dispersion coefficient as a function of $Pe\,B$ in the extensional flow. The leading-order asymptotic solution for small $Pe\,B$ in (40) is denoted by the dashed line. In the small $Pe\,B$ regime, the asymptotic solution agrees well with both the full numerical solution (circles) of the macrotransport equations and results obtained from BD simulations (squares). In the extensional flow, rotational dispersion is hindered and the dispersion coefficient vanishes in the large $Pe\,B$ limit. Because the extensional flow acts to align the particle orientation with the extensional axis, in the strong flow limit random reorientations are suppressed.
@@
where $\tilde{g}(\xi)=g(\phi)$. Noting that $\cos(2\phi_{0})=0$, and $\cos\left(2\phi_{0}+2\epsilon^{1/2}\xi\right)=-2\xi\sqrt{\epsilon}+\frac{4}{3}\xi^{3}\epsilon^{3/2}-\frac{4}{15}\xi^{5}\epsilon^{5/2}+O(\epsilon^{7/2})$, we expect an expansion of the form
@@
where the dots denote higher-order corrections. In figure 7(a), we compare the leading-order solution in (44) with the numerical solution of the full equation for $Pe\,B=20$. We note that indeed the leading-order solution in (44) approaches a delta function upon appropriate scaling in the limit $\epsilon\to 0$.
Figure 7: (a) Plots of numerical solution of $g$ (dashed) at $Pe\,B=20$ and the leading-order asymptotic solution (dotted) given in (44). (b) Plots of numerical solution of $b$ (dashed) at $Pe\,B=20$ and the leading-order asymptotic solution (dotted) given in (48).
We now proceed to analyze the displacement field, which in the boundary layer is expanded as
@@
In figure 7(b), we compare the leading-order solution of $b$ in (48) with the numerical solution of the full equation for $Pe\,B=20$.
@@ 4 Concluding remarks
In this paper, we have developed a generalized Taylor dispersion theory that describes the long-time rotational dispersion of a spheroidal particle in linear flows that is constrained to rotate in the velocity-gradient plane. As is standard for Taylor dispersion, the average drift and the effective dispersion are treated in a unified framework by leveraging a flux-averaging method in Fourier space. The results obtained from the continuum theory are corroborated by Brownian dynamics simulations of the orientational equation of motion. Using asymptotic analysis in the strong flow limit, we have showed that a simple shear enhances while an extensional flow hinders the long-time rotational dispersion. More specifically, we showed that $D^{\mathrm{eff}}/D_{R}=O(1)$ in simple shear and $D^{\mathrm{eff}}/D_{R}\to 0$ in extensional flow as $Pe\to\infty$ for a given nonzero $B$. These results reveal that the long-time rotational dispersion depends qualitatively on the characteristics of the background flow.
@@
While we focused on the long-time dispersion in steady flows, our generalized Taylor dispersion theory applies equally well to time-periodic flows such as oscillatory shear. In oscillatory shear, one needs to first obtain the long-time (time-dependent) solutions to the average and displacement fields. In addition to the cell average employed for steady linear flows, a time average over the oscillation period is needed to obtain the long-time transport coefficients.
@@
In linear flows as considered in this paper, the angular velocity of the particle due to the flow does not depend on the spatial position. This spatial uniformity allows us to consider the conservation equation in orientation space only. A difficulty would arise if one wishes to consider the rotational dispersion of a particle in flows in the presence of no-slip boundaries such as pressure-driven channel flows. In such a quadratic flow field, the angular velocity depends linearly on the transverse coordinate. Mathematically, the marginalization of $P(\bm{x},\bm{q},t)$ does not lead to a closed equation for $\Psi(\bm{q},t)$. In this case, one often needs to solve the full probability distribution $P(\bm{x},\bm{q},t)$ in order to calculate the long-time translational or rotational transport properties.
@@ Appendix A Asymptotic solution in simple shear
In this appendix, we discuss the asymptotic solution for nearly spherical particles in simple shear in 2D. Inserting the expansion (17) into (14), we obtain at steady state
@@
where $\Omega_{0}=-\frac{1}{2}Pe$, $\Omega_{1}=\frac{1}{2}Pe\cos(2\phi)$, and $n=0,1,\cdots$. In (59), $b_{k}=0$ for $k<0$. At $O(1)$, we have

### v2

@@ 3.2 Extensional flow
As another case study, we now consider an extensional flow where the angular velocity is $\bm{\Omega}=Pe\,B\cos(2\phi)\bm{e}_{z}$. The extensional flow tends to align the particle orientation with the extensional axis (see figure 5). The particle has a non-zero angular velocity when the orientation deviates from the extensional axis. As shown in figure 5(b), the direction of rotation depends on the particle orientation relative to the extensional axis.
@@
Figure 6: Plots of the effective dispersion coefficient as a function of $Pe\,B$ in the extensional flow. Circles are results obtained from the numerical solution of the Taylor dispersion theory, squares are results calculated from BD simulations and the dashed line denotes the leading-order asymptotic solution for small $Pe\,B$ given in (40).
In figure 6 we plot the effective dispersion coefficient as a function of $Pe\,B$ in the extensional flow. The leading-order asymptotic solution for small $Pe\,B$ in (40) is denoted by the dashed line. In the small $Pe\,B$ regime, the asymptotic solution agrees well with both the full numerical solution (circles) of the macrotransport equations and results obtained from BD simulations (squares). In the extensional flow, rotational dispersion is hindered and the dispersion coefficient vanishes in the large $Pe\,B$ limit. Because the extensional flow acts to align the particle orientation with the extensional axis, in the strong flow limit random reorientations are suppressed.
@@
where $\tilde{g}(\xi)=g(\phi)$. Noting that $\cos(2\phi_{0})=0$ and $\cos\left(2\phi_{0}+2\epsilon^{1/2}\xi\right)=-2\xi\sqrt{\epsilon}+\frac{4}{3}\xi^{3}\epsilon^{3/2}-\frac{4}{15}\xi^{5}\epsilon^{5/2}+O(\epsilon^{7/2})$, we expect an expansion of the form
@@
where the dots denote higher-order corrections. In figure 7(a) we compare the leading-order solution in (44) with the numerical solution of the full equation for $Pe\,B=20$. We note that indeed the leading-order solution in (44) approaches a delta function upon appropriate scaling in the limit $\epsilon\to 0$.
Figure 7: (a) Plots of the numerical solution of $g$ (dashed) at $Pe\,B=20$ and the leading-order asymptotic solution (dotted) given in (44). (b) Plots of the numerical solution of $b$ (dashed) at $Pe\,B=20$ and the leading-order asymptotic solution (dotted) given in (48).
We now proceed to analyse the displacement field, which in the boundary layer is expanded as
@@
In figure 7(b) we compare the leading-order solution of $b$ in (48) with the numerical solution of the full equation for $Pe\,B=20$.
@@
3.3 General linear flows
We now consider a general two-dimensional linear flow of the form $\bm{u}=\dot{\gamma}y\bm{e}_{x}+\alpha\dot{\gamma}x\bm{e}_{y}$, where $\alpha$ is a dimensionless parameter that controls the characteristics of the external flow. As a function of $\alpha$, the flow varies from purely rotational ($\alpha=-1$) to simple shear ($\alpha=0$) and extensional flow ($\alpha=1$). From (1), we obtain the angular velocity of the particle as
$\Omega\tau_{R}=\frac{1}{2}Pe\left[-1+\alpha+B\left(1+\alpha\right)\cos(2\phi)\right].$ (50)
Figure 8: Plots of the dispersion coefficient as a function of $\alpha$ for $Pe=100$ and $B=0.6$. Circles are results from numerical solutions of the macrotransport equations. Squares are results from BD simulations. The vertical line is given by the equation $\alpha=(1-B)/(1+B)$. The solid line is the asymptotic solution in equation (56).
To probe the effect of $\alpha$ on the long-time rotational dispersion coefficient, we solve the macrotransport equations (14) and (15) numerically using a Fourier collocation method for $Pe=100$. In figure 8 the numerical results are plotted as circles, and the results obtained from BD simulations are marked by squares. The dispersion coefficient exhibits a non-monotonic dependence on the flow parameter $\alpha$. The maximum of the dispersion coefficient is obtained when $\alpha=(1-B)/(1+B)$. For a general linear flow, the angular velocity is a combination of a constant part (purely rotational) and a varying part (extensional flow). For $\alpha=-1$, the flow is purely rotational, and we have $D^{\mathrm{eff}}=D_{R}$. As $\alpha$ increases, the varying part of the angular velocity gives enhanced dispersion. In the absence of the constant part ($\alpha=1$), particles becomes strongly aligned. As a result, the dispersion is suppressed (see section 3.2). To achieve maximum enhancement in dispersion, the two terms need to balance, which occurs when $1-\alpha=B(1+\alpha)$ or $\alpha=(1-B)/(1+B)$. In figure 8 this value of $\alpha$ is marked by a vertical line for $B=0.6$. On the right side of the vertical line, the extensional flow becomes dominant, and the dispersion coefficient decays to zero as $\alpha$ increases and approaches that of the extensional flow ($\alpha=1$).
In the large-$Pe$ limit, we consider a perturbation expansion in powers of $1/Pe$,
$g=g_{0}+\frac{1}{Pe}g_{1}+\cdots,$ (51a)
$b=b_{0}+\frac{1}{Pe}b_{1}+\cdots.$ (51b)
Following the notations and calculations in section 3.1, we obtain
$g_{0}=\frac{A_{1}}{1-\alpha-B(1+\alpha)\cos(2\phi)},$ (52a)
$g_{1}=\frac{4B(1+\alpha)A_{1}\sin(2\phi)}{\left[1-\alpha-B(1+\alpha)\cos(2\phi)\right]^{3}},$ (52b)
where $A_{1}=\sqrt{(1-\alpha)^{2}-B^{2}(1+\alpha)^{2}}$. We note that (52a) and (52b) are only valid in the region $1-\alpha>B(1+\alpha)$, which is shown as the shaded region in figure 9. At the boundary, $1-\alpha=B(1+\alpha)$, the solutions in (52a) and (52b) becomes singular at $\phi_{0}$, where $\cos(2\phi_{0})=1$ or $\phi_{0}\in\{0,\pi,2\pi\}$. One may similarly obtain the displacement field as
$\tilde{b}_{0}=A_{1}\frac{\arctan\left(\frac{1-\alpha+B(1+\alpha)}{A_{1}}\tan\phi\right)-\phi}{1-\alpha-B(1+\alpha)\cos(2\phi)},$ (53)
$\tilde{b}_{1}=\frac{C_{1}\left[\arctan\left(\frac{(1-\alpha+B(1+\alpha))}{A_{1}}\tan\phi\right)-\phi\right]+C_{2}}{\left[1-\alpha-B(1+\alpha)\cos(2\phi)\right]^{3}}+\frac{C_{3}}{1-\alpha-B(1+\alpha)\cos(2\phi)},$ (54)
where
$C_{1}=4B(1+\alpha)A_{1},\quad C_{2}=-3A_{1}^{2},$ (55a)
$C_{3}=\frac{B^{2}(1+\alpha)^{2}+(1-\alpha)\left(2(1-\alpha)+A_{1}\right)}{A_{1}^{2}}.$ (55b)
The leading-order dispersion coefficient can then be shown to be
$\frac{D_{0}^{\mathrm{eff}}}{D_{R}}=\frac{2(1-\alpha)^{2}+B^{2}(1+\alpha)^{2}}{2(1-\alpha)^{2}-2B^{2}(1+\alpha)^{2}}.$ (56)
For simple shear, $\alpha=0$, we recover equation (32). In figure 8 the solution in (56) for $B=0.6$ is plotted as a solid line. The asymptotic solution agrees well with both the numerical and BD results. As expected, the asymptotic solution deviates from the numerical or BD results as $\alpha$ approaches the singular boundary where $1-\alpha=B(1+\alpha)$. Furthermore, the solution of (56) become singular when $1-\alpha=B(1+\alpha)$.
In Appendix B, using asymptotic theory, we analyse the macrotransport equations and characterize the long-time dispersion behaviour for the case of $1-\alpha=B(1+\alpha)$ in the strong flow limit. When $1-\alpha=B(1+\alpha)$, our analysis reveals that there exist boundary layers of thickness $O(\epsilon^{1/3})$ at $\phi=n\pi$, $n\in\mathbb{Z}$, where $\epsilon=1/(Pe(1-\alpha))$. We show that the dispersion coefficient is $O(\epsilon^{-2/3})$ as $\epsilon\to 0$. For a fixed $\alpha$, this means that $D^{\mathrm{eff}}/D_{R}=O(Pe^{2/3})$.
Our results show that depending on the parameters $\alpha$ and $B$, the rotational dispersion coefficient exhibits distinctly different scaling in the large-$Pe$ limit. For $1-\alpha>B(1+\alpha)$, we have $D^{\mathrm{eff}}/D_{R}=O(1)$ as $Pe\to\infty$. On the curve where $1-\alpha=B(1+\alpha)$, $\alpha\neq 1$, $D^{\mathrm{eff}}/D_{R}=O(Pe^{2/3})$. Finally, for extensional flow $D^{\mathrm{eff}}/D_{R}\to 0$ as $Pe\to\infty$. We note that as a function of $\alpha$, the transition from the case of $1-\alpha=B(1+\alpha)$ to the vanishing dispersion at $\alpha=1$ occurs in a small region, which is shown by the rapid decay of $D^{\mathrm{eff}}$ in figure 8. Within this small region, a perturbation analysis for a small deviation from the curve $1-\alpha=B(1+\alpha)$ may be useful. Such an analysis is left for future considerations.
Figure 9: Plot of the parameter space $(\alpha,B)$. The solid line is given by $1-\alpha=B(1+\alpha)$. In the shaded region, we have $1-\alpha>B(1+\alpha)$.
@@ 4 Concluding remarks
In this paper we have developed a GTD theory that describes the long-time rotational dispersion of a spheroidal particle in linear flows that is constrained to rotate in the velocity-gradient plane. As is standard for Taylor dispersion, the average drift and the effective dispersion are treated in a unified framework by leveraging a flux-averaging method in Fourier space. The results obtained from the continuum theory are corroborated by BD simulations of the orientational equation of motion. Using asymptotic analysis in the strong flow limit, we have shown that a simple shear enhances while an extensional flow hinders the long-time rotational dispersion. More specifically, we showed that $D^{\mathrm{eff}}/D_{R}=O(1)$ in simple shear and $D^{\mathrm{eff}}/D_{R}\to 0$ in extensional flow as $Pe\to\infty$ for a given non-zero $B$. These results reveal that the long-time rotational dispersion depends qualitatively on the characteristics of the background flow.
@@
While we focused on the long-time dispersion in steady flows, our GTD theory applies equally well to time-periodic flows such as oscillatory shear. In oscillatory shear one needs to first obtain the long-time (time-dependent) solutions to the average and displacement fields. In addition to the cell average employed for steady linear flows, a time average over the oscillation period is needed to obtain the long-time transport coefficients.
@@
In linear flows as considered in this paper, the angular velocity of the particle due to the flow does not depend on the spatial position. This spatial uniformity allows us to consider the conservation equation in orientation space only. A difficulty would arise if one wishes to consider the rotational dispersion of a particle in flows in the presence of no-slip boundaries such as pressure-driven channel flows. In such a quadratic flow field, the angular velocity depends linearly on the transverse coordinate. Mathematically, the marginalization of $P(\bm{x},\bm{q},t)$ does not lead to a closed equation for $\Psi(\bm{q},t)$. In this case, one often needs to solve the full probability distribution $P(\bm{x},\bm{q},t)$ in order to calculate the long-time translational or rotational transport properties. As a concrete example, consider the rotational dynamics of a ‘point’ Brownian particle in a planar Poiseuille flow. The Smoluchowski equation can be written as
[…]
where $\bm{\Omega}$ depends on $y$, which is the transverse coordinate. Following the consideration given in section 2, one can show that the average field is governed by
$\frac{\partial}{\partial y}\left(-D_{T}\frac{\partial g}{\partial y}\right)+\frac{\partial}{\partial\phi}\left(\Omega(y)g-D_{R}\frac{\partial g}{\partial\phi}\right)=0.$ (58)
At the channel walls ($y=\pm H$), the no-flux condition, $-\partial g/(\partial y)=0$, is imposed. The normalization involves both spatial and orientational integrals, which is given by $\langle\int_{-H}^{H}g(y,\phi)\text{d}y\rangle=1$. Similarly, one may write down the governing equation for the displacement field.
The rotational Taylor dispersion theory can accommodate particles of different shapes such as chiral or ring-shaped particles (Singh et al., 2013). In the presence of hydrodynamic translation-rotation coupling, one again needs to consider the dynamics in both position and orientation space. In the case of general particle shapes, much like the case of a Brownian particle in Poiseuille flow, one performs the marginalization to obtain net orientational distributions after solving the average and displacement fields as functions of both $\bm{x}$ and $\bm{q}$.
The rotational Taylor dispersion theory is developed in two dimensions where the particle rotates about a single axis. To accommodate particles that rotate in three-dimensional space, one may consider the rotational dynamics about perpendicular axes. For a particular chosen axis of rotation, one can then utilize the current theory to characterize the long-time dynamics. We note that measuring rotational diffusion about perpendicular axes in three dimensions has been used in experiments (Rogers et al., 2012; Carrasco-Fadanelli et al., 2024). To conclude, we also note that similar to classical Taylor dispersion, our theory is only valid at long times.
@@ Appendix A Asymptotic solution in simple shear
In this appendix we discuss the asymptotic solution for nearly spherical particles in simple shear in two dimensions. Inserting the expansion (17) into (14), we obtain at steady state
@@
where $\Omega_{0}=-\frac{1}{2}Pe$, $\Omega_{1}=\frac{1}{2}Pe\cos(2\phi)$ and $n=0,1,\cdots$. In (68), $b_{k}=0$ for $k<0$. At $O(1)$, we have
@@
Appendix B Asymptotic analysis for $1-\alpha=B(1+\alpha)$ in the strong flow limit
We develop an asymptotic solution in the large-$Pe$ limit for a general linear flow under the constraint that $1-\alpha=B(1+\alpha)$. Under this condition, we have $\Omega\tau_{R}=-Pe(1-\alpha)\sin^{2}\phi$. For convenience, we define $\epsilon=1/(Pe(1-\alpha))$ and consider the behaviour of the macrotransport equations when $\epsilon\ll 1$. We note that $\alpha\neq 1$ is assumed.
The dimensionless average field at steady state, $g=g(\phi)$, is governed by
$\frac{\text{d}}{\text{d}\phi}(\sin^{2}\phi g)+\epsilon\frac{\text{d}^{2}g}{\text{d}\phi^{2}}=0.$ (74)
Noting that $g(\phi+\pi)=g(\phi)$, we consider $\phi\in[-\pi/2,\pi/2]$. The normalization is given by
$\frac{1}{\pi}\int_{-\pi/2}^{\pi/2}g(\phi)\text{d}\phi=1.$ (75)
An asymptotic solution to equation (74) subject to the normalization condition (75) has been developed by Brenner (1970). Since the solution to $g$ is required for the consideration of the displacement field $b$, we first repeat the calculations of Brenner (1970) for $g$ and then proceed to consider $b$.
Integrating (74) once, we obtain
$\sin^{2}\phi g+\epsilon\frac{\text{d}g}{\text{d}\phi}=C^{\prime}(\epsilon),$ (76)
where we remark that $C^{\prime}$ only depends on $\epsilon$. In the bulk, we have $\sin^{2}\phi\;g\sim C^{\prime}$ or $g\sim C^{\prime}/\sin^{2}\phi$. This bulk result breaks down when $\sin^{2}\phi\,g\sim\epsilon\,\text{d}g/(\text{d}\phi)$ or $\phi=O(\epsilon^{1/3})$, which suggests the existence of a boundary layer with thickness $O(\epsilon^{1/3})$ at $\phi=0$. In the boundary layer, from (76), we obtain $g\sim C^{\prime}\epsilon^{-2/3}$. Using the normalization condition (75), one can show that $C^{\prime}=O(\epsilon^{1/3})$. Following Brenner (1970), we define $C^{\prime}=C\epsilon^{1/3}$, where $C=O(1)$. In the bulk, we define $g(\phi)=C\epsilon^{1/3}H(\phi;\epsilon)$. Introducing the stretched coordinate $s=\phi/\epsilon^{1/3}$, we define $g(s)=C\epsilon^{-1/3}Q(s;\epsilon)$ in the boundary layer. Expanding $H$ and $Q$, we write
$H=H_{0}+\epsilon H_{1}+\cdots,\quad Q=Q_{0}+\epsilon^{2/3}Q_{1}+\cdots,$ (77)
where $H_{0}=1/\sin^{2}\phi$ and $Q_{0}$ is governed by
$s^{2}Q_{0}(s)+\frac{\text{d}Q_{0}}{\text{d}s}=1.$ (78)
The solution to $Q_{0}$ is given by
$Q_{0}(s)=e^{-s^{3}/3}\int_{-\infty}^{s}e^{r^{3}/3}\text{d}r.$ (79)
To leading order, the normalization condition reduces to
$C\int_{-\infty}^{+\infty}Q_{0}(s)\text{d}s=\pi,$ (80)
which allows us to obtain
$C=\frac{3^{5/6}\pi}{6\sqrt{3}\left(\Gamma\left(4/3\right)\right)^{2}+2^{4/3}\sqrt{\pi}\Gamma\left(7/6\right)}=\frac{3^{1/3}\pi}{[\Gamma(1/3)]^{2}},$ (81)
where $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\text{d}t$ is the gamma function and $C\approx 0.6313$. We can construct a leading-order composite solution to $g$ as $C\epsilon^{-1/3}Q_{0}(\phi/\epsilon^{1/3})$. From this, we obtain the leading-order drift as $\Omega^{\mathrm{eff}}\tau_{R}=-C\epsilon^{-2/3}+o(\epsilon^{-2/3})$. In figure 10(a) we compare the asymptotic result with the results from numerical solutions of the macrotransport equations for $B=1$ and $\alpha=0$ in the large-$Pe$ regime.
The displacement field is governed by
$-\frac{\text{d}}{\text{d}\phi}\left(\sin^{2}\phi\,b\right)-\epsilon\frac{\text{d}^{2}b}{\text{d}\phi^{2}}=2\epsilon\frac{\text{d}g}{\text{d}\phi}+\epsilon\left(\Omega^{*\mathrm{eff}}-\Omega^{*}\right)g,$ (82)
where $\Omega^{*\mathrm{eff}}=\Omega^{\mathrm{eff}}\tau_{R}$ and $\Omega^{*}=\Omega\tau_{R}$. Noting that the dominant term on the right-hand side outside the boundary layer is $\epsilon\Omega^{*}g=O(\epsilon^{1/3})$, we have $b=O(\epsilon^{1/3})$ in the bulk. In the boundary layer, a dominant balance reveals that $b=O(\epsilon^{-1/3})$. While the resulting equations are not analytically tractable, the above analysis allows us to show that $D^{\mathrm{eff}}/D_{R}=O(\epsilon^{-2/3})$ as $\epsilon\to 0$. In figure 10 we present the dispersion coefficient obtained from numerical solutions of the macrotransport equations for $B=1$ and $\alpha=0$ in the large-$Pe$ regime. The solid line has the form $D^{\mathrm{eff}}/D_{R}=\tilde{C}\epsilon^{-2/3}$, where $\tilde{C}$ is a numerical constant that fits the data closely.
Figure 10: (a) Plots of the numerical and asymptotic solutions to the average angular drift velocity as a function of $1/\epsilon$ for $B=1,\alpha=0$. (b) Plots of the numerical solutions to the non-dimensional effective long-time dispersion coefficient as a function of $1/\epsilon$ for $B=1,\alpha=0$. The solid line is given by $D^{\mathrm{eff}}/D_{R}=\tilde{C}\epsilon^{-2/3}$, where $\tilde{C}$ is obtained from numerical data.